OpenStudy (anonymous):

Solve: 2x-20 -------- =2x 3 1) -5 2)-3 3)2 4) 13

OpenStudy (anonymous):

\[\frac{ 2x - 20 }{ 3 } = 2x\]

OpenStudy (anonymous):

We're going to plug in and work out each one.

OpenStudy (anonymous):

I'll start.

OpenStudy (anonymous):

\[\frac{ 2(-5) - 20 }{ 3} = 2(-5)\] Work this one out.